A transparent plastic has been developed. When light travels through it, it only travels at 3 × 107 m/s. What will be the refractive index of this new plastic?


1
Expert's answer
2021-07-19T16:03:59-0400

"n = \frac{c}{v}"

"v = 3*10^7 \frac{m}{s}"

"c\approx 3*10^8\frac{m}{s}"

"n = \frac {3*10^8}{3*10^7}= 10"


"\text{Answer: the refractive index n = 10}"